4.         Sales are planned for a department at $2,345,698 for the Fall season. The seasonal turnover objective is 1.6.

Calculate the BOM stock for the months of August, September and December if sales are planned at $199,384, $342,643, and $703,709 respectively and the basic stock method is used.

August BOM Stock:

September BOM Stock:

December BOM Stock:

please show your work, thank you
Answer #1
We would use the following formulas
i. Basic Stock= Average Stock for the Season - Average Monthly Sales for the Season
where
Average Monthly Sales for the Season = Total Planned Sales for the Season / Number of months in the Season
Average Stock for the Season = Total Planned Sales for the Season / Estimated Inventory Turnover Rate for the Season
Beginning of the Month (BOM) stock = Planned Monthly Sales + Basic Stock
Fall Season is generally for 3 months
thus Average Monthly Sales for the Season = 2345698/3 = 781899.33
Average Stock for the Season = 2345698/1.6 = 1466061.25
thus Basic Stock = 1466061.25 - 781899.33 = 684161.92
Thus August BOM Stock : Planned Monthly Sales (199384) + Basic Stock (684162 rounding off)
: 199384 + 684162 = 883546
Thus September BOM Stock : Planned Monthly Sales (342643) + Basic Stock (684162 rounding off)
: 342643 + 684162 = 1026805
Thus December BOM Stock : Planned Monthly Sales (703709) + Basic Stock (684162 rounding off)
: 703709 + 684162 = 1387871